Prove that the equation of the chord joining the points P(ct, c/t) and Q(cT, c/T) on
the rectangular hyperbola xy = c
2
is x + tT y = c(t + T). M is the midpoint of P Q
and P Q meets the x-axis at N. Prove that OM = MN, where O is the origin
1
Expert's answer
2020-03-17T14:13:06-0400

P (ct,c/t) and Q (cT,c/T)

Slope of line joining PQ is="\dfrac{(c/T – c/t)}{(cT – ct)}= – 1/Tt"

 Equation of the line using slope-point form is "(y\u2013ct)=(\u20131\/cTt)(x\u2013c\/t)"


Simplifying we get "x+tTy=c(t+T)"


M is the midpoint of PQ. By midpoint formula,


"M\u2261(c(t+T)\/2,c(1\/t+1\/T)\/2)"


"i.e.M(c(t+T)\/2,c(t+T)\/2tT)"

By distance formula,

"OM=\\sqrt{(\\frac{(2c(t+T)\u200b\u20130)}2)^2+(\\frac{c(t+T)}{2tT}\u200b\u20130)^2}"

Simplify to get OM ="[c(t + T)\sqrt{(1 + T^2t^2)] / 2}] …...(1)"

Put y = 0 in the equation of PQ for the coordinates of N.

"N\u2261(c(t+T),0)"


Similarly MN ="\sqrt{{(\dfrac{c(t + T)}2 – c(t + T)})^2 + {{(\dfrac{c(t + T)}{2tT} – 0}})^2}"


Simplifying we get MN="[c(t+T)(1+T^2t^2)\/2\n\u200b]\u2026...(2)"

From (1) & (2),

OM = MN